Hello,
I did ask this question: The ages of Peter, Margaret and Jack are consecutive odd numbers. The sum of the ages of Peter and Margaret equals Jack’s age before 5 years. Find the ages of Peter, Margaret and Jack.
I read the solution but I don't know how you get the consecutive numbers: x, x + 2 and x + 4. Can you please explain to me why?
Thank you.

Answer by richwmiller(17219)

consecutive odd numbers would be numbers like 5 7 9
5
7 is 5+2
9 is 5+4
let's call the first number x
x=5
x+2=5+2=7
x+4=5+4=9